OpenStudy (anonymous):

Which object exerts more gravitational force, object A with a mass of 25 grams and a density of 2 g/cm3 or object B with a mass of 55 grams and a density of 0.5 g/cm3?

OpenStudy (anonymous):

the first one, as 25 * 2 = 50 , and 55 * 0.5 = 27.5.

OpenStudy (carlsmith):

This may seem like a daft question, sorry, but I thought gravitation was a product of mass alone, I didn't think density was a factor?? I'd be grateful if someone could put me straight on that.

OpenStudy (anonymous):

If both objects are in vacuum, the only factor affects the gravitation force is their mass (drag doesn't exist). Weight of A = 0.025 * g (g = \[9.8 m/s^{2}\]), and weight of B = 0.055 *g. A is lighter, meaning gravitation force on A is less that B.

OpenStudy (anonymous):

"an object exerting gravitational force" this is definitely a weird question. what we call gravitational force, is the force of attraction between two bodies and it is defined as: F = G *{m1* m2}*{r^2} if you talk about which object is SUBJECT to earth gravitational force, then F=mg just as mlutfi in the previous message explained it. in all the cases it has nothing to do with density